A bag contains 8 red jelly beans and 8 green jelly beans. How many ways can 5 jelly beans be withdrawn from the bag with the number of green ones being less than 3?

Answer by robertb(5830)

The 3 possibilities:
3 red, 2 green;
4 red, 1 green;
5 red, 0 green.
The first case: 8C3*8C2 = 56*28 = 1,568.
The secind case: 8C4 * 8C1 = 70*8 = 560.
The third case: 8C5 * 8C0 = 56*1 = 56.
Adding these figures arising from the mutually exclusive cases, we get 1,568 + 560 + 56 = 2,184 ways.